Dr. David Bryan Carpenter Flat 8, Broughton House 28/30 Bullar Road Bitterne Park Southampton, S018 1GS

December 6, 1995

The Visa Service, 2 Northdown Street, Kings Cross, London, N1 9BG

Dear Visa Service,

I am moving to the United States for two years, starting January 1996, to work at an American university. I need to obtain a J-1 Exchange Visitor visa.

I hold a full United Kingdom passport, and have of an IAP 66 form from my future employer.

Please could you send me the appropriate application forms and any relevant information. I enclose a deposit of £10 in the form of a cheque made payable to “The Visa Service”.

Yours sincerely,